OpenStudy (anonymous):

Integrate (x^2+4)/(x+2) Leave answer as integral in terms of u.

myininaya (myininaya):

have you tried diving the integrand?

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{x^2+4}{x+2}dx}\) is same as, \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{(x+2)(x-2)}{x+2}dx}\) I am not sure why you need a "u" here....

OpenStudy (solomonzelman):

just cancel things out, and integrate term by term....

myininaya (myininaya):

x^2-4=(x+2)(x-2) x^2+4=(x+2i)(x-2i)

OpenStudy (solomonzelman):

ooops

OpenStudy (solomonzelman):

yes, the second I put that, I was like nope... sorry \(\large\color{slate}{\displaystyle\int\limits_{~}^{~}\frac{x^2+4}{x+2}dx}\)

OpenStudy (solomonzelman):

try to make the denominator a signle term (using u substitution)

OpenStudy (solomonzelman):

*single term

OpenStudy (solomonzelman):

what would your "u" be?

OpenStudy (anonymous):

u=x+2

myininaya (myininaya):

oh wait are we not suppose to integrate? it says leave integrals in terms of u... I guess if that is so let u=x+2 then du=dx and blah blah

OpenStudy (anonymous):

Would you rewrite as (x^2+4)(x+2)^-1

myininaya (myininaya):

put still it doesn't say put the integrand in a form that is integrateble

myininaya (myininaya):

weird question

myininaya (myininaya):

but (not put)

OpenStudy (solomonzelman):

(As I just showed I suck, lol, but) You can would need to integrate, I guess.

OpenStudy (solomonzelman):

I mean just need.

myininaya (myininaya):

well it says "leave answer as integral in terms of u"

myininaya (myininaya):

it doesn't say integrate

myininaya (myininaya):

it just says to convert \[\int\limits_{}^{}f(x) dx \text{ to } \int\limits_{}^{}g(u) du\]

myininaya (myininaya):

and if that is so then you can use any substitution (useless or not)

ganeshie8 (ganeshie8):

guess these divisions always have shortcuts \[\frac{x^2+4}{x+2} = \dfrac{(x^2-4)+8}{x+2} = x-2 + \frac{8}{x+2}\] i feel "answer" refers to integrate, i also feel the OP is hiding some details of actual problem hmm

OpenStudy (solomonzelman):

magic zero, I sometimes refer to what ganeshie just did. I like that.

OpenStudy (solomonzelman):

Idk if you have to integrate or not, but again, I think you would still need to. This is calculus afterall, not just an algebraic substitution technique.

ganeshie8 (ganeshie8):

@phanta1 say something.. what does \(u\) refer to here ? if psble please take a screenshot of actual problem and attach

myininaya (myininaya):

lol I just noticed my first posting says diving instead dividing